OpenStudy (anonymous):

A textbook is 23cm wide. What is this in meters?

OpenStudy (shubhamsrg):

23 cm = 23* 10^(-5) km also, 1 kn = 10^3 m hope that helps! ;)

OpenStudy (anonymous):

100 cm = 1 meter. If you start with 23 cm and multiply in a way that makes the units cancel, you'll have your answer. \[23cm \times \frac{ 1m }{ 100cm } =0.23m\] The math is just 23/100 but the units are what tell you what numbers to multiply or divide by.